Find the y-intercept.
–x + 3y = 15
Thanks.

Answer by rapaljer(4671)

The y-intercept is always where x=0. So substitute x=0 into the equation, and you get
-x+3y=15
-0+3y=15
3y=15
y=5

R^2 at SCC